### v1

@@ 3.2 Bias identification with heteroscedastic noise
In comparison, the three approaches converge to a similarly wrong value of Young’s modulus in the fitted model due to the inability of the computational system to reflect the observations with the additional deformation. The optimal value for $E$ such that either the MSE or the $L^{2}$ loss between observations and predictions are minimized is around 14.8 GPa, which is approximately reached by each of the approaches. As both MSE and $L^{2}$-optimum values are relatively close, it was expected that both the OGP and the unbiased method converge to a similar value. KOH’s approach tends to those same values, as these optimum parameters would provide the best-fitting bias distribution for the chosen GP prior. Adding the noise as a tunable parameter in the approach without bias does not improve the result nor quantify the real noise at the output. Contrarily, adding the bias term as in either approach allows the correction of the bias and quantifies the noise reliably. In this case, there is no apparent benefit of using OGP over KOH despite its larger computational times (see Table 4), as the chosen prior for KOH leads to the same optimum. As the bias term is directly applied at the output of the system, the value of $E$ does not reflect the one in the generative model and therefore could not be used for derived QoI that depend on it.
@@
The key feature highlighted by these results is the capacity of the bias term to represent a deviation of a biased model from reality. The noise in the approach without bias term cannot be quantified appropriately due to the existing discrepancy, while KOH can reliably correct the prescribed displacement and infer the noise produced by the variations in the material. This physical noise component of the observations is included as part of the model bias, which agrees with the assumption that the model’s Young’s modulus employed in the digital twin is unique and deterministic. The inability of the model to represent stochastic responses is compensated by a noise-enabled bias with a modified kernel. Without this bias term, the model response variance remains unknown and the model predictions are not able to represent the reality of the system. It is shown that a heteroscedastic noise kernel can be used in the bias to this end in a non-intrusive manner, without modifying the model or the inference schemes. Samples from this posterior predictive distribution can be used in the context of a digital twin to provide information on points where no observation is available or to further derive QoI based on such predictions. The stochastic nature of the bias-corrected predictions provides a quantification of the uncertainty in the observations produced by the variability of the real system, and not only due to the indeterminacy of the model parameters. For example, large variations in the bias could indicate a deficient testing setup, or in this case, the inclusion of loads not considered in the model.
@@ 3.3.1 Case definition
The objective of this case is the inclusion of additional information on the environmental conditions in the bias identification applied to a real use case. Consequently, the final use case illustrates their implementation on a simplified model of an actual bridge: the Nibelungenbrücke in Worms, Germany, displayed in Figure 10. The research initiative "Schwerpunktprogramm 2388: SPP Hundert Plus", within which this work is included, aims to monitor and eventually prolong the remaining life of the bridge through novel digital approaches, in particular the implementation of digital twins46. Therefore, a model using a simplified version of Nibelungenbrücke’s geometry is developed, paying special attention to its use in such a digital twin. The objective of this use case is the evaluation of the model bias approaches in a digital twin setting and the inclusion of additional variables, therefore the modelling of the bridge has been extremely simplified and would not necessarily coincide with an accurate representation of the real one. It must be particularly noted that the spans of the real bridge should be represented as cantilever structures attached at the center of the spans and with the pilots as supports. However, the simulation of a span will be represented by a simply-supported structure of the full span length for illustrative purposes. This choice has no impact on the conclusions extracted from this benchmark with respect to the application of biased approaches. The work presented in this paper will use simulation models for both generating the virtual observation data and predicting results. Measurements from the bridge are expected in the future, which will allow further testing of the methodology under real sensor observations. However, the validation with real data remains out of the scope of this paper, and this benchmark reflects the real system with many simplifying assumptions.
@@
[…]
@@
The computational model to be tuned will be the simulation to generate the influence lines and the parameter of interest that must be inferred will be the Young’s modulus used in the simulation. An additional temperature effect is added only for the generation of the dataset, introducing a discrepancy between the generated data and the computational model. During the time that the truck would pass over the bridge, it is reasonable to assume a change in the temperature of the top surface of the deck49. This can be, for example, due to the incidence of the sun over it. At the start of the experiment, we support a constant stationary temperature over the whole bridge. The top part of the deck will increase its temperature due to the transient effect of the heat on the surface, while the bottom part will remain unaffected due to the short timescale that prevents fast heat propagation. This situation generates a linear temperature gradient along the cross-section of the deck, which implies further thermal strains that will be identified by the sensor. These strains are introduced in the linear elastic model with a thermal expansion coefficient $\alpha=10^{-5}$. A diagram of the experimental set-up is represented in Figure 10. In practice, the difference in temperature between the start and end of the measurements on the top of the deck is measured, supposing a linear evolution of it. In this case, we run 6 simulations with temperature differences $\Delta T$ between start and end of [1e-2, 2.8e-2, 4.6e-2, 6.4e-2, 8.2e-2, 1e-1] K. This corresponds to a change of between 5 and 15 K in a period of 12 h, which is in line with expected values in similar applications50. The observed influence lines for this experiment are represented in Figure 11. The base case with no temperature difference has been represented in the figure as a reference, but will not be fed as data to the model for its calibration. Notably, the deformation is greater than zero at the end of the simulation when the temperature gradient is introduced. This effect cannot be replicated by the original model, as the only external load applied on the system comes from the passing truck, which at the end of the measurements is no longer over the bridge. Therefore, all the fitted models will predict zero deformation at the end of the simulation. No additional noise is introduced in the system, although the same methodology of the previous benchmark could be applied.
@@ 3.3.2 Models without temperature consideration
In comparison, the imposition of the orthogonality condition onto the bias GP kernel leads the distribution of $E$ to converge towards the MSE. It must be noted that the correction generated by the bias from OGP only fits the observations in the case of an optimal value of $E$. The orthogonality condition restricts -through a modification of the covariance matrix- the function space that can represent the bias correction. In this case, as the derivative of the computational model with respect to the Young’s modulus does not change much in the studied range of $E$, the orthogonal kernel remains almost constant for every evaluation. In particular, the functions generated by the bias GP will resemble the case where $E$ is optimal, the MSE of the observations. Additionally, the orthogonal kernel translates into an impossibility of the bias GP to be trained to fit residuals that would be improved by a modification of the latent parameter. In this case, it means that, for an arbitrary $E$, the bias will be able to correctly compensate for the discrepancy at the end of the experiment (which can only be explained through the bias) but not the values when the load is at the center of the span (which can be largely altered by a modification in the value of $E$). These diminished effects of the bias on some parts of the domain imply larger covariance values due the inability to cover the observed residuals with members of the imposed function space. The increased loss of smoothness due to high variance for values of $E$ far from 57.8 GPa can be observed in Figure 15. Larger variations in the derivative of the computational model with respect to $E$ would lead to more dissimilar realizations of the bias response mean. Nevertheless, the convergence to the optimum $E$ is achieved thanks to diminishing covariances obtained by the fulfilment of the orthogonality condition in values closer to the MSE. The mean of the bias response corresponds with the residuals for that value in the whole domain of application.
@@ 3.3.3 Models with temperature consideration
In this case, the bias GP will depend on the temperature difference as well. Its main attraction is to include information that may be relevant to the process without modifying the computational model itself. This extension of the domain of interest through the bias allows weighting in the influence of the temperature difference in the inference process while providing insight into the potential model deficiencies. Therefore, the bias GP will be trained on pairs of the form $(x,\Delta T)$, where $x$ is the current load position and $\Delta T$ is the associated end temperature difference for a given observation. This bias can consequently be represented by a response surface with a 2D input and a 1D output. Each series of observations is part of the same experiment with an associated $\Delta T$ measured at the end of it. The intermediate temperatures are not available, therefore all the observations will share the final $\Delta T$ as their coordinate in the extended domain. The training of the bias term and the inference procedures are implemented following the algorithms from Section 2. In this case, an additional scaling of the domain and the residuals is applied to improve the fitting of the bias GP. The GP inputs are scaled such that the original domain $(x,\Delta T):[0,104]\times[1e-2,1e-1]$ is transformed to $[0,1]\times[0,1]$. The residuals and derivatives are multiplied by a factor of 100 during the training to increase numerical stability. The originally small residuals produced by the displacements lead to values of the variance close to machine precision after scaling the time domain, which leads to numerical instabilities of the GP training. To compensate this effect, the scaling factor in the residuals is chosen in the same order of magnitude as the original time length. The same previous kernels are applied for KOH and OGP, extending them to accept 2D inputs.
@@ 4 Conclusion
In this paper, we have proposed solutions to some of the main challenges for the reliable implementation of simulation-based digital twins of bridges. A Bayesian inference framework is often considered the best option for the calibration of the model parameters. However, we observed that classical approaches led to overconfident models whose response does not replicate the observations and whose behaviour does not correspond to the physical system. This renders them unsuitable for their use in digital twins of critical infrastructures such as bridges, where fidelity, robustness and accuracy are key to guarantee their safety. The root of this inadequacy stems from the inability of computational models to recreate the infinitely complex reality to its fullest extent. We proposed the explicit introduction of a model bias term in the formulation and calibration of the digital twin to mitigate such an effect. Two different approaches were compared with the classical inference methodology, one with a modularized version of Kennedy and O’Hagan’s framework and one based on OGPs. Through three relevant use cases, we proved that biased approaches did not only compensate for the model bias but also provide new insight into model deficiencies. Additionally, we introduced two novel extensions to the model bias identification schemes to correctly quantify noise in the observations from unknown physical sources and to make use of complementary information.
@@
However, one of the main challenges of the presented methodologies is the inability to assure the physicality of the inferred parameters due to the addition of the corrective bias term at the output of the predictions. The model itself is therefore not directly modified, and the corrections cannot be utilized for the calculation of different QoIs than the ones used for the calibration. Nevertheless, the generated bias and the corrected models allowed us to better understand the deficiencies present in the current model in a quantitative manner. Additionally, the evaluation of the presented frameworks under real measurements is required for validation of the exposed approaches in structures as-built. The complexity of sensor observations and the variation in environmental and operational conditions cannot be replicated with simulated data, therefore the evaluation of a definitive assessment of their performance is only possible under such conditions. Furthermore, the integration of these methodologies in a unified digital twin of a bridge presents additional challenges related to accessibility, ownership, connectivity and cybersecurity that should not be understated.
@@ Acknowledgments
This work was funded through the project "C07 - Data driven model adaptation for identifying stochastic digital twins of bridges" from the Priority Program (SPP) 2388/1 "Hundred plus" of the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG).

### v2

@@ 3.2 Bias identification with heteroscedastic noise
In comparison, the three approaches converge to a similarly wrong value of Young’s modulus in the fitted model due to the inability of the computational system to reflect the observations with the additional deformation. The optimal value for $E$ such that either the MSE or the $L^{2}$ loss between observations and predictions are minimized is around 14.8 GPa, which is approximately reached by each of the approaches. As both MSE and $L^{2}$-optimum values are relatively close, it was expected that both the OGP and the unbiased method converge to a similar value. KOH’s approach tends to those same values, as these optimum parameters would provide the best-fitting bias distribution for the chosen GP prior. Adding the noise as a tunable parameter in the approach without bias does not improve the result nor quantify the real noise at the output. Contrarily, adding the bias term as in either approach allows the correction of the bias and quantifies the noise reliably. In this case, there is no apparent benefit of using OGP over KOH despite its larger computational times (see Table 6), as the chosen prior for KOH leads to the same optimum. As the bias term is directly applied at the output of the system, the value of $E$ does not reflect the one in the generative model and therefore could not be used for derived quantities that depend on it.
@@
The key feature highlighted by these results is the capacity of the bias term to represent a deviation of a biased model from reality. The noise in the approach without bias term cannot be quantified appropriately due to the existing discrepancy, while KOH can reliably correct the prescribed displacement and infer the noise produced by the variations in the material. This physical noise component of the observations is included as part of the model bias, which agrees with the assumption that the model’s Young’s modulus employed in the digital twin is unique and deterministic. The inability of the model to represent stochastic responses is compensated by a noise-enabled bias with a modified kernel. Without this bias term, the model response variance remains unknown and the model predictions are not able to represent the reality of the system. It is shown that a heteroscedastic noise kernel can be used in the bias to this end in a non-intrusive manner, without modifying the model or the inference schemes. Samples from this posterior predictive distribution can be used in the context of a digital twin to provide information on points in $X$ where no observation is available. The stochastic nature of the bias-corrected predictions provides a quantification of the uncertainty in the observations produced by the variability of the real system, and not only due to the indeterminacy of the model parameters. For example, large variations in the bias could indicate a deficient testing setup, or in this case, the inclusion of loads not considered in the model.
@@ 3.3.1 Case definition
The objective of this case is the inclusion of additional information on the environmental conditions in the bias identification applied to a real use case. Consequently, the final use case illustrates their implementation on a simplified model of an actual bridge: the Nibelungenbrücke in Worms, Germany, displayed in Figure 10. The research initiative "Schwerpunktprogramm 2388: SPP Hundert Plus", within which this work is included, aims to monitor and eventually prolong the remaining life of the bridge through novel digital approaches, in particular the implementation of digital twins62. Therefore, a model using a simplified version of Nibelungenbrücke’s geometry is developed. The objective of this use case is the evaluation of the model bias approaches in a digital twin setting and the inclusion of additional variables, therefore the modelling of the bridge has been extremely simplified and would not necessarily coincide with an accurate representation of the real one. It must be particularly noted that the spans of the real bridge should be represented as cantilever structures attached at the center of the spans and with the pilots as supports. However, the simulation of a span will be represented by a simply-supported structure of the full span length for illustrative purposes. This choice has no impact on the conclusions extracted from this benchmark with respect to the application of biased approaches. The work presented in this paper will use simulation models for both generating the virtual observation data and predicting results. Measurements from the bridge are expected in the future, which will allow further testing of the methodology under real sensor observations. However, the validation with real data remains out of the scope of this paper, and this benchmark reflects the real system with many simplifying assumptions. Table 7 presents a summary of the formulation for this case.
@@
Element Formulation
Generator model assumptions Load line-test of linear elastic simplified bridge model that depends on the Young’s modulus $E$, the position $x$ of a known test truck load of 10 t and the temperature difference $\Delta T$ between start and end of the experiment, and outputs the vertical displacement $u$ at $x=47$ m.
Generator model $u^{g}(x,\Delta T,E)=g(x,\Delta T,E)+\epsilon$
Computational model assumptions Load line-test of linear elastic simplified bridge model that depends on the Young’s modulus $E$ and the position $x$ of a known test truck load of 10 t, and outputs the vertical displacement $u$ at $x=47$ m.
Computational model without bias term $u^{f}(x,E)=f(x,E)+\epsilon$
Computational model with bias term (without temperature) $u^{h}(x,E)=u^{f}(x,\theta)+b(x)=f(x,E)+b(x)+\epsilon$
Computational model with bias term (with temperature) $u^{h}(x,\Delta T,E)=u^{f}(x,\theta)+b(x,\Delta T)=f(x,E)+b(x,\Delta T)+\epsilon$
Discrepancy source Additional $u$ induced by $\Delta T$
Latent parameter and prior $E,\leavevmode\nobreak\ p(E)\sim\mathcal{LN}(35.8\text{ GPa},8.9\text{ GPa})$
Domain of interest Load position $x\in X=[0\text{ m},104\text{ m}]$$\times$ $\Delta T\in\Theta=[0.01\text{ K},0.1\text{ K}]$
Measurements’ domain $(\hat{x},\hat{\Delta T})\in(\hat{X}\times\hat{\Theta})=\left([0\text{ m},0.01\text{ K}],[104\text{ m},0.1\text{ K}]\right)$, observations at $\Delta T=$[0.01, 0.028, 0.046, 0.064, 0.082, 0.1] K for $x$ at intervals of 1 m.
Noise $\epsilon\sim\mathcal{N}(0,\sqrt{1\times 10^{-12}})$
Table 7: Case description for Section 3.3 (bridge model). The terms $f(x,E)$ and $g(x,E)$ represent the output of the FE model of the system under the assumptions of the computational model and the generator, respectively.
The first span of the bridge is chosen for the simulation model. It is represented as a solid 3D Finite Element (FE) box-girder model with variable cross-section. The dimensions of the model correspond to those indicated in available drawings63 and databases64, with a span length of 95,185 m, span width varying between 14 m at the centre and 14.64 at the pilots, deck thickness of 0.25 m, cross-section height at pilots of 6.5 m and of 2.5 m at the middle of the span. The mesh is a tetrahedral 3D solid unstructured mesh with linear Lagrangian elements, with characteristic length between 0.2 and 2.0 m. The material is modelled using a purely linear elastic constitutive law with a Young’s modulus of 40 GPa, a Poisson’s ratio of 0.2 and a density of 2350kg/m3. Simply supported boundary conditions are considered at the end surfaces, where the bottom edge of the supports is fixed in displacements, as represented in Figure 9. The resulting model has 3530 nodes, which is equivalent to 10580 Degrees of Freedom (DoFs) and 20669 elements. A more complex model is possible, but it would not enhance in any way the conclusions from this study.
@@
The computational model to be tuned will be the simulation to generate the influence lines and the parameter of interest that must be inferred will be the Young’s modulus used in the simulation. An additional temperature effect is added only for the generation of the dataset, introducing a discrepancy between the generated data and the computational model. During the time that the truck would pass over the bridge, it is reasonable to assume a change in the temperature of the top surface of the deck65. This can be, for example, due to the incidence of the sun over it. At the start of the experiment, we support a constant stationary temperature over the whole bridge. The top part of the deck will increase its temperature due to the transient effect of the heat on the surface, while the bottom part will remain unaffected due to the short timescale that prevents fast heat propagation. This situation generates a linear temperature gradient along the cross-section of the deck, which implies further thermal strains that will be identified by the sensor. These strains are introduced in the linear elastic model with a thermal expansion coefficient $\alpha=10^{-5}$. A diagram of the experimental set-up is represented in Figure 10. In practice, the difference in temperature between the start and end of the measurements on the top of the deck is measured, supposing a linear evolution of it. In this case, we run 6 simulations with temperature differences $\Delta T$ between start and end of [0.01, 0.028, 0.046, 0.064, 0.082, 0.1] K. This corresponds to a change of between 5 and 15 K in a period of 12 h, which is in line with expected values in similar applications66. The observed influence lines for this experiment are represented in Figure 11. The base case with no temperature difference has been represented in the figure as a reference, but will not be fed as data to the model for its calibration. Notably, the deformation is greater than zero at the end of the simulation when the temperature gradient is introduced. This effect cannot be replicated by the original model, as the only external load applied on the system comes from the passing truck, which at the end of the measurements is no longer over the bridge. Therefore, all the fitted models will predict zero deformation at the end of the simulation. No additional noise is introduced in the system except for a negligible value added for numerical stability, although the same methodology of the previous benchmark could be applied.
@@ 3.3.2 Models without temperature consideration
In comparison, the imposition of the orthogonality condition onto the bias GP kernel leads the distribution of $E$ to converge towards the MSE. It must be noted that the correction generated by the bias from OGP only fits the observations in the case of an optimal value of $E$. The orthogonality condition restricts, through a modification of the covariance matrix, the function space that can represent the bias correction. In this case, as the derivative of the computational model with respect to the Young’s modulus does not change much in the studied range of $E$, the orthogonal kernel remains almost constant for every evaluation. In particular, the functions generated by the bias GP will resemble the case where $E$ is optimal, the MSE of the observations. Additionally, the orthogonal kernel translates into an impossibility of the bias GP to be trained to fit residuals that would be improved by a modification of the latent parameter.For an arbitrary $E$, the bias will be able to correctly compensate for the discrepancy at the end of the experiment (which can only be explained through the bias) but not the values when the load is at the center of the span (which can be largely altered by a modification in the value of $E$). These diminished effects of the bias on some parts of the domain imply larger covariance values due the inability to cover the observed residuals with members of the imposed function space. The increased loss of smoothness due to high variance for values of $E$ far from 57.8 GPa can be observed in Figure 15. Larger variations in the derivative of the computational model with respect to $E$ would lead to more dissimilar realizations of the bias response mean. Nevertheless, the convergence to the optimum $E$ is achieved thanks to diminishing covariances obtained by the fulfilment of the orthogonality condition in values closer to the MSE. The mean of the bias response corresponds with the residuals for that value in the whole domain of application.
@@ 3.3.3 Models with temperature consideration
A possible alternative to cover the temperature variations in the system is to include the noise standard deviation $\sigma_{\epsilon}$ in the latent parameters instead of prescribing it, as it was presented in Section 3.2 for the application without bias term. However, this does not take advantage of the temperature information nor increases the model’s predictive capacity.
Alternatively, we make the bias GP depend on the temperature difference as well $b(x)\to b(x,\Delta T)$. Its main interest resides in the inclusion of information that may be relevant to the process without modifying the computational model itself. This extension of the domain of interest through the bias allows weighting in the influence of the temperature difference in the inference process while providing insight into the potential model deficiencies. Therefore, the bias GP will be trained on pairs of the form $(x,\Delta T)$, where $x$ is the current load position and $\Delta T$ is the associated end temperature difference for a given observation. This bias can consequently be represented by a response surface with a 2D input and a 1D output. Each series of observations is part of the same experiment with an associated $\Delta T$ measured at the end of it. The intermediate temperatures are not available, therefore all the observations will share the final $\Delta T$ as their coordinate in the extended domain. The training of the bias term and the inference procedures are implemented following the algorithms from Section 2. In this case, an additional scaling of the domain and the residuals is applied to improve the fitting of the bias GP. The GP inputs are scaled such that the original domain $(x,\Delta T):[0,104]\times[0.01,0.1]$ is transformed to $[0,1]\times[0,1]$. The residuals and derivatives are multiplied by a factor of 100 during the training to increase numerical stability. The originally small residuals produced by the displacements lead to values of the variance close to machine precision after scaling the time domain, which leads to numerical instabilities of the GP training. To compensate this effect, the scaling factor in the residuals is chosen in the same order of magnitude as the original time length. The same previous kernels are applied for KOH and OGP, extending them to accept 2D inputs.
@@ 4 Conclusion
In this paper, we have proposed solutions to some of the main challenges for the reliable implementation of simulation-based digital twins of bridges. A Bayesian inference framework is often considered the best option for the calibration of the model parameters. However, we observed that classical approaches led to overconfident models whose response does not replicate the observations and whose behaviour does not correspond to the physical system. This renders them unsuitable for their use in digital twins of critical infrastructures such as bridges, where fidelity, robustness and accuracy are key to guarantee their safety. The root of this inadequacy stems from the inability of computational models to recreate the infinitely complex reality to its fullest extent. We proposed the explicit introduction of a model bias term in the formulation and calibration of the digital twin to mitigate such an effect. Two different approaches were compared with the classical inference methodology, one with a modularized version of Kennedy and O’Hagan’s framework and one based on OGPs. Through three relevant use cases, we proved that biased approaches did not only compensate for the model bias but also provide new insight into model deficiencies. Additionally, we introduced two novel extensions to the model bias identification schemes to correctly quantify noise in the observations from unknown physical sources and to make use of complementary information. The bias identified through these extensions can be used as a bases for guiding potential model improvements.
@@
However, one of the main challenges of the presented methodologies is the inability to assure the physicality of the inferred parameters due to the addition of the corrective bias term at the output of the predictions. The model itself is therefore not directly modified, and the corrections cannot be utilized for the calculation of different quantities than the ones used for the calibration. Nevertheless, the generated bias and the corrected models allowed us to better understand the deficiencies present in the current model in a quantitative manner. Additionally, the evaluation of the presented frameworks under real measurements is required for validation of the exposed approaches in structures as-built. The complexity of sensor observations and the variation in environmental and operational conditions cannot be replicated with simulated data, therefore the evaluation of a definitive assessment of their performance is only possible under such conditions. Furthermore, the integration of these methodologies in a unified digital twin of a bridge presents additional challenges related to accessibility, ownership, connectivity and cybersecurity that should not be understated.
@@ Acknowledgments
This work was funded through the project “C07 - Data driven model adaptation for identifying stochastic digital twins of bridges” from the Priority Program (SPP) 2388/1 “Hundred plus” of the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG).